# Resource-Theoretic Hierarchy of Contextuality for General Probabilistic Theories

### v1

@@ 4.2 Contextuality of composite systems
When studying a resource theory, it is often also important to specify how resource objects are composed to obtain a joint resource. As discussed in Section 3.1, GPT systems do not have a canonical tensor product. However, any sensible choice of a tensor product $\mathbin{\tilde{\otimes}}$ should give rise to a symmetric monoidal category with exact simulations as morphisms [58, Section 5.2]. Since the transformations in our resource theory of GPT-contextuality are not merely simulations of GPT systems, but also involve additional classical systems, it takes a bit more work to show that they (and the free transformations) are closed under parallel composition. We prove this in the rest of this section for the case of free operations.
@@
Assuming that there is a unique composite of any GPT system with a classical one given by the canonical composite of Definition 5 (see [63] for an alternative composition of classical systems) and that $\mathbin{\tilde{\otimes}}$ is associative, the codomain can be expressed as
@@ 4.3 Quantifying contextuality via the classical excess
In order to construct a specific contextuality monotone, we will make use of the excess function from Definition 20. Of particular interest is $\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})$, the $\mathsf{A}$-excess within the countable classical GPT system (Example 4). For brevity, we call it the classical excess of $\mathsf{A}$. A crucial ingredient in the following proof is that $\Delta_{\mathbb{N}}$ is isomorphic to $\Delta_{\mathbb{N}}\otimes\Delta_{n}$ for any finite $n$.
@@
is a direct consequence of Proposition 19 and the definition of excess.
For the converse inequality, note that any $\epsilon$-embedding $(\phi,\psi):\mathsf{A}\to\Delta_{\mathbb{N}}$ can be extended to an $\epsilon$-embedding
$(\phi\otimes\mathrm{id},\psi\otimes\mathrm{id})\;:\;\mathsf{A}\otimes\Delta_{n}\to\Delta_{\mathbb{N}}\otimes\Delta_{n},$ (58)
@@
Furthermore, there is a 00-embedding $\Delta_{\mathbb{N}}\otimes\Delta_{n}\to\Delta_{\mathbb{N}}$ given by the linear map
@@
where $(\omega_{i})_{i\in\mathbb{N}}$ is an element of $\ell^{1}$, $\delta_{j}$ is a basis vector in $\mathbb{R}^{n}$, and $\delta_{j|i}$ returns $1$ if $j$ divides $i$ and 00 otherwise. Using basically the same linear map for effects, merely extended to $\ell^{\infty}\otimes\mathbb{R}^{n}\to\ell^{\infty}$, gives the required embedding. Consequently, we have
@@
so that by inequality (59) and Proposition 21 we get
@@
By Definition 23, Corollary 22 (specialized to $\mathsf{C}=\Delta_{\mathbb{N}}$), and Lemma 25 respectively, we have
$\mathsf{A}\preceq\mathsf{B}$ $\iff\quad\exists\,n\;\text{ such that }\;\mathsf{A}\hookrightarrow\mathsf{B}\otimes\Delta_{n},$ (64)
@@
An important feature of the monotone $\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})$ is that it can sometimes be computed by considering the $\mathsf{A}$-excess within $\Delta_{k}$ for a finite $k$. This happens whenever its value $\varepsilon(\mathsf{A},\Delta_{m})$ remains constant for all classical systems with $m$ larger than $k$. For instance, the noncontextual GPT system corresponding to the toy bit in Spekkens toy theory [66] is embedabble within $\Delta_{4}$ with vanishing error, and hence for any larger $m$ its excess within $\Delta_{m}$ has to remain 00.
@@
Having introduced the hierarchy of contextuality as well as a new contextuality measure in the form of the classical excess, we now build a measure of generalised contextuality based on the parity oblivious multiplexing (POM) [10] game. The latter is a protocol that is powered by preparation contextuality and has raised significant attention in recent years [67, 68, 69, 70, 71, 72, 73, 74].
@@
Note that the string ${x}$ is assumed to be uniformly distributed, so that the left-hand side of Equation 65 can be interpreted (up to a factor of $2^{n-1}$) as the average state sent by Alice given that the characters of ${x}$ add up to 00 modulo $2$ and similarly for the right-hand side.
@@
It is shown in [10, Theorem 2] that if a system $\mathsf{A}$ is noncontextual, then the optimal (i.e. maximal) success probability for the $n$-bit POM, denoted by $\mathfrak{p}_{n}$, is upper bounded by
@@
Let us consider any choice of states $\omega_{x}$ (for $x\in\{0,1\}^{n}$) which obey the parity obliviousness constraint (65). Since we have $\mathsf{A}\hookrightarrow\mathsf{B}$, for any choice of effects $e_{y}$ and states $\omega_{x}$ on $\mathsf{A}$ obeying the parity obliviousness constraint, the states $\Gamma(\omega_{x})$ on $\mathsf{B}$ also obey it by the linearity of the single-valued state simulation $\Gamma$. Moreover, the empirical adequacy (Equation 26) of the exact simulation implies $e_{y}\cdot\omega_{x}=\Gamma(e_{y})\cdot\Gamma(\omega_{x})$. Thus, any strategy on $\mathsf{A}$ can be implemented on $\mathsf{B}$ with equal success probability. This implies $\mathfrak{p}_{n}(\mathsf{A})\leq\mathfrak{p}_{n}(\mathsf{B})$. ∎
@@
By the desiderata 1 and 2, a noisy bit $\Delta_{2}^{\alpha}$ and a bit $\Delta_{2}$ must satisfy $\Delta_{2}^{\alpha}\preceq\Delta_{2}$ and $\Delta_{2}\preceq\Delta_{2}^{\alpha}$ since they are both noncontextual. The optimal strategy for POM given a (noisy) bit consists of encoding the first bit $x_{1}$ of the string $x$ in its extremal states $\delta_{1}$ and $\delta_{2}$ [10, Lemma 1]. With probability $1/n$ Bob is asked to guess this specific bit, which he can do successfully with probability $(1-\alpha)$. With probability $(n-1)/{n}$ Bob is asked to guess any of the remaining bits from $x$, for which he can do no better than a uniformly random guess, i.e. the success probability is $1/2$. Hence, the optimal success probability is
@@
which is strictly smaller than $\mathfrak{p}_{n}(\Delta_{2})=(n+1)/2n$ whenever the noise parameter $\alpha$ is greater than 00. This means that $\Delta_{2}\preceq\Delta_{2}^{\alpha}$ does not imply $\mathfrak{p}_{n}(\Delta_{2}^{\alpha})\geq\mathfrak{p}_{n}(\Delta_{2})$. ∎
We can see that the POM success probability cannot serve as a monotone because the POM game does not allow the use of classical systems for free. This can be remedied by the so-called generalized yield construction [65, Section 3.1].
@@
Let us assume $\mathsf{A}\preceq\mathsf{B}$, which implies that there is an $m\in\mathbb{N}$ such that there is a 00-embedding $\mathsf{A}\otimes\Delta_{d}\to\mathsf{B}\otimes\Delta_{d\cdot m}$ for every $d\in\mathbb{N}$. By Lemma 28, we obtain $\mathfrak{p}_{n}(\mathsf{A}\otimes\Delta_{d})\leq\mathfrak{p}_{n}(\mathsf{B}\otimes\Delta_{d\cdot m})$ for every $d\in\mathbb{N}$, which yields the inequality below:
@@
using Definition 30. ∎
This proof is an instance of the proof of the generalized yield construction in [26, Theorem 4.21], so that $\mathfrak{p}_{n}{\textrm{-}}\mathrm{yield}$ can be meaningfully interpreted as a (generalized) yield monotone.
@@ 5.1 Contextuality and information erasure
The fine-tuning problem of contextuality.
Contextuality of a theory implies the existence of distinctions at the ontological level which are not present at the operational level. If a contextual ontological model truly describes the physical reality underlying the observed behaviours predicted by the theory, then there are operationally indistinguishable behaviours that have distinct ontological origins. In other words, such operational equivalences would result from a fine-tuning of the corresponding distinct ontological representations [4]. How do these distinctions disappear between the ontological and operational descriptions of the physical system, though? The presence of such fine-tunings provides a conspiratorial connotation to the realist explanation of the theory and we believe that it requires an explanation. In this section, we explore the possibility that the fine-tuning associated with contextuality can be explained as emergent from a yet undiscovered physical mechanism that supplements the description provided by the ontological model.
@@
Explaining the origin of fine-tunings of this kind by searching for new physical mechanisms dates back to Valentini’s veriant of Bohmian mechanics [75]. There, he introduces a notion of quantum equilibrium as the reason why superluminal signaling does not manifest in quantum theory, despite the nonlocality of its underlying ontological model. This picture predicts that outside of the quantum equilibrium, it is possible to observe faster than light signaling. Therefore, the fine-tuned nature of no-signaling in Bohmian mechanics is explained just as an emergent feature of the quantum equilibrium and it is not universally valid. We cannot avoid noticing how radical such explanations of fine-tunings rooted in undiscovered physical mechanisms are. They imply that an established physical principle, such as the principle of no-signaling, is violated at the fundamental level. In the case of contextuality, the physical mechanism explaining the emergence of the operational equivalences would entail the existence of measurements that can distinguish behaviours that are deemed indistinguishable by quantum theory.
@@
In Valentini’s work the quantum equilibration process is responsible for the emergence of no-signalling — the fine-tuned feature associated with nonlocality. What hypothetical physical mechanism could be responsible for the emergence of operational equivalences — the fine-tuned feature associated with contextuality?
@@
It would have to be a process that involves a kind of information erasure. The information erased is the information about distinctions at the ontological (e.g. fundamental) level, which cannot be stored in systems of the operational (e.g. effective) theory that lacks these distinctions. By Landauer’s principle, we can then associate an increase in entropy between the fundamental and effective levels. Such a process of information erasure would not only provide an explanation for the problematic fine-tuning associated with contextuality but would also be associated to a potentially detectable heat dissipation. This heat would signify that, indeed, there are distinctions at the fundamental level which are not present at the effective level. One could even hypothesise that the information erasure is a physical process occuring over time. That is, during the preparation of a quantum system there may be a timescale before which the system is described by the fundamental (and noncontextual) theory. At longer timescales, once the erasure has occured, the system can only be described by the effective (contextual) theory.
Ontological model as a fundamental theory.
The above account treats an ontological model as specifying a yet unfalsified fundamental theory, in accordance with [32]. This departs from the standard use of ontological models to study contextuality [2]. In the latter, the ontological distinctions (which are not present in the operational theory) are in principle indistinguishable. On the other hand, the fundamental theory we posit here contains no such requirement. In particular, the in-principle-indistinguishability of its distinctions would also mean that there is no entropy increase and no heat resulting from erasure to be detected. Examples of effective theories arising from more fundamental ones include thermodynamics, which emerges from statistical mechanics via coarse-graining, and classical information theory, which emerges from quantum information theory via decoherence. The difference between the two interpretations of ontological models does not prevent us from using the approach of [32] to provide an explanation of generalized contextuality as defined in [2]. If the fine-tuning associated with generalized contextuality is explained through a process information erasure, then the problematic aspect of contextuality in quantum theory disappears. Instead, one is then led to search for an in principle accessible more fundamental theory, from which quantum theory emerges.
Resource theory of GPT-contextuality and physical erasure processes.
In our work, both the effective and fundamental theory are associated with GPT systems. Consider $\mathsf{A}$ as the effective description of a physical system and $\mathsf{B}$ the fundamental one. A simulation $\mathsf{A}\to\mathsf{B}$ can be then interpreted as specifying, for each behaviour of the effective theory, all behaviours of the fundamental theory that would give rise to it. For example, if the effective description arises from the fundamental one due to a physical erasure process given by some many-to-one function $f:\mathsf{B}\to\mathsf{A}$, then the associated simulation would be given by preimages under $f$. This point of view is consonant with the use of GPT simulations in [32] where they were introduced. However, not every simulation can be associated with a physical map (in the opposite direction) in such a way. Here, a physical map $\mathsf{B}\to\mathsf{A}$ would be a linear map $f:V_{\mathsf{B}}\to V_{\mathsf{A}}$ satisfying $f(\Omega_{\mathsf{B}})\subseteq\Omega_{\mathsf{A}}$, $f^{*}(E_{\mathsf{A}})\subseteq E_{\mathsf{B}}$ and $f^{*}(e_{\mathsf{A}})\cdot\omega_{\mathsf{B}}=e_{\mathsf{A}}\cdot f(\omega_{\mathsf{B}})$. For instance, while the Spekkens toy bit [66] has a noncontextual ontological model (and thus a univalent simulation) defined on the tetrahedron $\Delta_{4}$, there is no physical map from $\Delta_{4}$ to the toy bit. Some ontological models do correspond to physical maps, though. For any GPT system $\mathsf{A}$ with $n$ extremal states $\omega_{i}$ there exists a Holevo-Beltrametti-Bugasjki (HBB) model of $\mathsf{A}$ by $\Delta_{n}$, as shown in Example 16. The HBB simulation corresponds to a linear map $M:\mathbb{R}^{n}\to V_{\mathsf{A}}$ whose action on the extremal states of $\Delta_{n}$ is given by: $\delta_{i}\mapsto\omega_{i}$. See Appendix A for a proof that $M$ is a physical map.
@@
By analyzing physical maps like the one of the HBB simulation we can understand why information erasure in the context of ontological models and contextuality cannot be the same as the standard thermodynamic one. While the latter can be understood as a coarse-graining operation, the physical maps associated to contextual ontological models are not of this kind. To see this, consider a contextual GPT system $\mathsf{A}$ and a classical system $\Delta_{d}$ such that a physical map $f:\Delta_{d}\to\mathsf{A}$ exists. Since $\mathsf{A}$ is contextual, by Lemma 11 of [32], there is no coarse-graining (i.e. idempotent) process on $\Delta_{d}$ whose image is $\mathsf{A}$. Hence, the physical GPT map $f$ cannot be understood simply as a coarse-graining operation, such as the erasure that occurs between fundamental and effective levels in statistical mechanics. A more concrete perspective is offered by the HBB map $M$ which explicitly models the preparation of the system $\mathsf{A}$ via the classical control $\Delta_{n}$. The classical degree of freedom is used to select the extremal state of $\mathsf{A}$ to be prepared. Thus, while $M$ can be used to model a physical process in the laboratory (e.g. the preparation of $\mathsf{A}$ conditioned on a classical variable), it does not correspond to a simple coarse-graining process applied to the classical degree of freedom.

### v2

@@ 4.2 Contextuality of composite systems
When studying a resource theory, it is often also important to specify how resource objects are composed to obtain a joint resource. As discussed in Section˜3.1, GPT systems do not have a canonical tensor product. However, any sensible choice of a tensor product $\mathbin{\tilde{\otimes}}$ should give rise to a symmetric monoidal category with exact simulations as morphisms [66, Section 5.2]. Since the transformations in our resource theory of GPT-contextuality are not merely simulations of GPT systems, but also involve additional classical systems, it takes a bit more work to show that they (and the free transformations) are closed under parallel composition. We prove this in the rest of this section for the case of free operations.
@@
Assuming that there is a unique composite of any GPT system with a classical one given by the canonical composite of Definition˜5 (see [5, 6] for a derivation of the uniqueness of the composition with a classical system under the assumption of local tomography and [26] for an alternative composition of classical systems without this assumption) and that $\mathbin{\tilde{\otimes}}$ is associative, the codomain can be expressed as
@@ 4.3 Quantifying contextuality via the classical excess
In order to construct a specific contextuality monotone, we will make use of the excess function from Definition˜11. Of particular interest is $\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})$, the $\mathsf{A}$-excess within the countable classical GPT system (Example˜4). For brevity, we call it the classical excess of $\mathsf{A}$. A crucial ingredient in the following proof is that $\Delta_{\mathbb{N}}$ is isomorphic to $\Delta_{\mathbb{N}}\otimes\Delta_{n}$ for any finite $n$.
@@
is a direct consequence of Proposition˜10 and the definition of excess.
For the converse inequality, note that any univalent $\epsilon$-simulation $(\Gamma,\Theta):\mathsf{A}\to\Delta_{\mathbb{N}}$ can be canonically extended to a univalent $\epsilon$-simulation
$\bigl(\Gamma\otimes\mathrm{id},\Theta\otimes\mathrm{id}\bigr)\;:\;\mathsf{A}\otimes\Delta_{n}\to\Delta_{\mathbb{N}}\otimes\Delta_{n},$ (43)
@@
Furthermore, there is (an exact) univalent simulation $\Delta_{\mathbb{N}}\otimes\Delta_{n}\to\Delta_{\mathbb{N}}$ given on states by the linear map
@@
where $(\omega_{i})_{i\in\mathbb{N}}$ is an element of $\ell^{1}$, $\delta_{j}$ is a basis vector in $\mathbb{R}^{n}$, and $\delta_{j|i}$ returns $1$ if $j$ divides $i$ and 0 otherwise. Using basically the same linear map for effects, merely extended to $\ell^{\infty}\otimes\mathbb{R}^{n}\to\ell^{\infty}$, gives the required simulation. Consequently, we have
@@
so that, by Inequality (44) and Proposition˜12, we get
@@
By Definition˜14, Corollary˜13 (specialized to $\mathsf{C}=\Delta_{\mathbb{N}}$), and Lemma˜16 respectively, we have
[…]
@@
An important feature of the monotone $\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})$ is that it can sometimes be computed by considering the $\mathsf{A}$-excess within $\Delta_{k}$ for a finite $k$. This happens whenever its value $\varepsilon(\mathsf{A},\Delta_{m})$ remains constant for all classical systems with $m$ larger than $k$. For instance, the noncontextual GPT system corresponding to the toy bit in Spekkens toy theory [81] is embedabble within $\Delta_{4}$ with vanishing error, and hence for any larger $m$ its excess within $\Delta_{m}$ has to remain 0.
@@
Having introduced the hierarchy of contextuality as well as a new contextuality measure in the form of the classical excess, we now build a measure of generalized contextuality based on the parity oblivious multiplexing (POM) [80] game. The latter is a protocol that is powered by preparation contextuality and has raised significant attention in recent years [7, 21, 32, 69, 4, 86, 17, 50].
@@
Note that the string x{x} is assumed to be uniformly distributed, so that the left-hand side of Equation˜50 can be interpreted (up to a factor of $2^{n-1}$) as the average state sent by Alice given that the characters of x{x} add up to 0 modulo $2$ and similarly for the right-hand side.
@@
It is shown in [80, Theorem 2] that if a system $\mathsf{A}$ is noncontextual, then the optimal (i.e. maximal) success probability for the $n$-bit POM, denoted by $\mathfrak{p}_{n}$, is upper bounded by
@@
Let us consider any choice of states $\omega_{x}$ (for $x\in\{0,1\}^{n}$) which obey the parity obliviousness constraint (50). Since we have $\mathsf{A}\hookrightarrow\mathsf{B}$, for any choice of effects $e_{y}$ and states $\omega_{x}$ on $\mathsf{A}$ obeying the parity obliviousness constraint, the states $\Gamma(\omega_{x})$ on $\mathsf{B}$ also obey it by the linearity of the single-valued state simulation $\Gamma$. Moreover, the empirical adequacy (Equation˜18) of the exact simulation implies $e_{y}\cdot\omega_{x}=\Gamma(e_{y})\cdot\Gamma(\omega_{x})$. Thus, any strategy on $\mathsf{A}$ can be implemented on $\mathsf{B}$ with equal success probability. This implies $\mathfrak{p}_{n}(\mathsf{A})\leq\mathfrak{p}_{n}(\mathsf{B})$. ∎
@@
By the desiderata 1 and 2, a noisy bit $\Delta_{2}^{\alpha}$ and a bit $\Delta_{2}$ must satisfy $\Delta_{2}^{\alpha}\preceq\Delta_{2}$ and $\Delta_{2}\preceq\Delta_{2}^{\alpha}$ since they are both noncontextual. The optimal strategy for POM given a (noisy) bit consists of encoding the first bit $x_{1}$ of the string $x$ in its extremal states $\delta_{1}$ and $\delta_{2}$ [80, Lemma 1]. With probability $1/n$ Bob is asked to guess this specific bit, which he can do successfully with probability $(1-\alpha)$. With probability (n1)/n(n-1)/{n} Bob is asked to guess any of the remaining bits from $x$, for which he can do no better than a uniformly random guess, i.e. the success probability is $1/2$. Hence, the optimal success probability is
@@
which is strictly smaller than $\mathfrak{p}_{n}(\Delta_{2})=(n+1)/2n$ whenever the noise parameter $\alpha$ is greater than 0. This means that $\Delta_{2}\preceq\Delta_{2}^{\alpha}$ does not imply $\mathfrak{p}_{n}(\Delta_{2}^{\alpha})\geq\mathfrak{p}_{n}(\Delta_{2})$. ∎
We can see that the POM success probability cannot serve as a monotone because the POM game does not allow the use of classical systems for free. This can be remedied by the so-called generalized yield construction [35, Section 3.1].
@@
Let us assume $\mathsf{A}\preceq\mathsf{B}$, which implies that there is an $m\in\mathbb{N}$ such that there is a univalent simulation $\mathsf{A}\otimes\Delta_{d}\to\mathsf{B}\otimes\Delta_{d\cdot m}$ for every $d\in\mathbb{N}$. By Lemma 19, we obtain $\mathfrak{p}_{n}(\mathsf{A}\otimes\Delta_{d})\leq\mathfrak{p}_{n}(\mathsf{B}\otimes\Delta_{d\cdot m})$ for every $d\in\mathbb{N}$, which yields the inequality below:
@@
using Definition˜21. ∎
This proof is an instance of the proof of the generalized yield construction in [36, Theorem 4.21], so that $\mathfrak{p}_{n}{\textrm{-}}\mathrm{yield}$ can be meaningfully interpreted as a (generalized) yield monotone.
@@ 5 Discussion
This section is divided into two subsections. In the first, we discuss whether our resource-theoretic perspective on contextuality could offer an avenue to understand contextuality as an expression of information erasure. In the second, we discuss how our work relates to the main studies on which it builds, as well as to the resource theory of generalized contextuality developed in [28].
@@ 5.1 Contextuality and information erasure
The resource-theoretic perspective on contextuality [80, 41, 74, 69, 56, 57, 90, 30, 67, 28, 65] highlights the following aspect: the existence of a noncontextual model (i.e. noncontextuality) of a physical system imposes constraints on the type of information that this system can carry. This is in essence why one can show that its violation (i.e. contextuality) can act as a resource for information-theoretic tasks. Our resource theory of GPT-contextuality aims to give a more precise description of the nature and properties of this information-theoretic manifestation of contextuality as compared to previous works that often focus merely on the distinction between contextual and noncontextual systems.
From a more foundational perspective, a contextual ontological model gives a classical description of an operational theory that postulates ontological distinctions that are in principle inaccessible. There are two types of responses to the phenomenon of contextuality (e.g. that of quantum systems). One is to reject ontological models as the appropriate framework to explain the operational features of the theory. The other is to accept the possibility that a contextual ontological model provides a feasible realist explanation of the empirical predictions of the theory.
The latter, however, begs for an explanation of the fine-tuning associated to it [20]. How can we understand that ontological distinctions disappear at the operational level? One could explain the presence of fine-tuning implied by contextuality if there was a physical process responsible for this disappearance. It would map the ontological theory to the operational theory, effectively erasing the extra information that is supposed to be inaccessible at the level of the operational theory. We refer to this hypothetical process, associated to any ontological model, as the revelation map, for short.
One could then ask: can revelation maps explain the detailed information-theoretic properties of contextuality? To tackle this question, we may take our resource theory of GPT-contextuality as an expression of the information-theoretic properties of contextuality. In this sense, the question then becomes: can the resource theory of GPT-contextuality be alternatively interpreted as a resource theory of erasure for revelation maps? We cannot currently answer this question, since a resource theory of erasure for revelation maps has not yet been developed. Nevertheless, we believe the question is worthwhile and we outline a few relevant considerations below.
@@
The presence of fine-tunings in a contextual ontological model gives a conspiratorial connotation to the realist explanation of the theory it provides. One attempt to justify the presence of such fine-tunings, by searching for new physical mechanisms, dates back to Valentini’s variant of Bohmian mechanics [87]. There, he introduces a notion of quantum equilibrium as the reason why superluminal signaling does not manifest in quantum theory, despite the nonlocality of its underlying ontological model. This picture predicts that outside of the quantum equilibrium, it is possible to observe faster than light signaling. Therefore, the fine-tuned nature of no-signaling in Bohmian mechanics is explained just as an emergent feature of the quantum equilibrium and it is not universally valid. We cannot avoid noticing how radical such explanations of fine-tunings rooted in undiscovered physical mechanisms are. They imply that an established physical principle, such as the principle of no-signaling, is violated at the fundamental level. In the case of contextuality, the physical mechanism explaining the emergence of the operational equivalences would entail the existence of measurements that can distinguish behaviours that are deemed indistinguishable by quantum theory.
@@
In Valentini’s work, the quantum equilibration process is responsible for the emergence of no-signalling — the fine-tuned feature associated with nonlocality. What hypothetical physical mechanism could be responsible for the emergence of operational equivalences — the fine-tuned feature associated with contextuality?
@@
It would have to be a process that involves a kind of information erasure. The information erased is the information about distinctions at the ontological (e.g. fundamental) level, which cannot be stored in systems of the operational (e.g. effective) theory that lacks these distinctions. By Landauer’s principle, we can then associate an increase in entropy between the fundamental and effective levels. Such a process of information erasure would not only provide an explanation for the problematic fine-tuning associated with contextuality but would also be associated to a potentially detectable heat dissipation. This heat would signify that, indeed, there are distinctions at the fundamental level which are not present at the effective level. One could even hypothesise that the information erasure is a physical process occurring over time. That is, during the preparation of a quantum system there may be a timescale before which the system is described by the fundamental (and noncontextual) theory. At longer timescales, once the erasure has occurred, the system can only be described by the effective (contextual) theory.
Empirical nature of the fundamental theory.
In the above hypothetical account of Bohmian mechanics, the theory of Bohmian particles is seen as a more fundamental one than quantum theory.
In order for a similar explanation of the fine-tuning associated with contextuality to be testable, the more fundamental theory must be in principle accessible. Only then would we expect entropy increase and heat transfer as a result of the erasure in the revelation map. This is in accordance with the view presented by Müller and Garner in [64], but it is not in opposition to the standard use of ontological models to study contextuality [79].
Indeed, if one tries to find evidence for the physical erasure process we hypothesize here and fails, this gives more credence to the common assumption that any ontological distinctions, which are not present in the operational theory, are in principle indistinguishable. On the other hand, if such evidence is found, the detailed study of contextuality in quantum theory provides invaluable clues towards finding the more fundamental theory, from which quantum theory emerges.
Ontological models as physical erasure processes.
In case the fundamental theory in question is classical, one is justified in using standard ontological models to describe how the effective theory is simulated by the fundamental one. Namely, an ontological model of (the operational theory associated to) a GPT system $\mathsf{A}$ (Section˜3.2) is equivalent to a simulation of $\mathsf{A}$ by a classical system $\Delta_{n}$. The classical system $\Delta_{n}$ is interpreted as the fundamental system and $\mathsf{A}$ as the effective description thereof. A simulation $\mathsf{A}\to\Delta_{n}$ specifies all behaviours of the fundamental theory $\Delta_{n}$ that are compatible with a given behaviour of the effective theory $\mathsf{A}$.
Whether we can interpret $\mathsf{A}$ as a description of $\Delta_{n}$ that emerges as a result of a revelation map from $\Delta_{n}$ to $\mathsf{A}$ depends on the kind of revelation maps that we consider. For example, if it supposed to be a pair of partial functions (one for states, one for effects) $f:\Delta_{n}\to\mathsf{A}$, then the associated simulation would be given by preimages under $f$. We could obtain an arbitrary simulation in this way.
However, many simulations cannot be obtained like this if we further require the revelation map $f$ to be a physical GPT transformation (Definition˜23). Namely, we say that a simulation $(\Gamma,\Theta)$ of $\mathsf{A}$ by $\Delta_{n}$ admits of a physical realisation if there exists a linear map $M:\Delta_{n}\to\mathsf{A}$, such that $\Gamma(\omega)\subseteq M^{-1}(\omega)$ for all $\omega\in\Omega_{\mathsf{A}}$ and $M^{*}(e)\in\Theta(e)$ for all $e\in E_{\mathsf{A}}$. See Appendix˜A for more details.
For instance, while there is a noncontextual ontological model of the Spekkens toy bit [81] by two classical bits, i.e. a univalent simulation by $\Delta_{4}$, there cannot exist a physical realisation of this simulation, as we prove in Appendix˜A.
Nevertheless, there are ontological models with physical realisations. For any GPT system $\mathsf{A}$ with a finite set $\{\omega_{i}\}_{i=1}^{n}$ of convexly extremal states, we can define the Holevo-Beltrametti-Bugajski (HBB) model of $\mathsf{A}$ (Example˜9), which may be contextual. In Lemma˜27, we show that every HBB model, viewed as a simulation of $\mathsf{A}$ by a classical GPT system $\Delta_{n}$, admits of a physical realisation $M:\Delta_{n}\to\mathsf{A}$. At the level of states, $M$ maps a given extremal state of the simplex $\Delta_{n}$ to the corresponding extremal state $\omega_{i}$ of $\Omega_{\mathsf{A}}$.
@@
By analyzing physical GPT transformations like $M$ above, we can understand why information erasure in the context of revelation maps cannot be the same as the erasure in the context of thermodynamics. While the latter can be understood as a coarse-graining operation, physical maps associated to contextual ontological models are not necessarily of this kind.
To see this, consider a contextual ontological model of a GPT system $\mathsf{A}$, which is given by preimages under a physical GPT transformation $f:\Delta_{d}\to\mathsf{A}$. Since $\mathsf{A}$ is contextual, there is no coarse-graining (i.e. idempotent) process on $\Delta_{d}$ whose image is $\mathsf{A}$ [64, Lemma 11]. Hence, the physical GPT transformation $f$ cannot be understood simply as a coarse-graining operation, such as the erasure that occurs between fundamental and effective levels in statistical mechanics.
A more concrete perspective is offered by the HBB map $M$ which explicitly models the preparation of the system $\mathsf{A}$ via the classical control $\Delta_{n}$. The classical degree of freedom is used to select the extremal state of $\mathsf{A}$ to be prepared. Thus, while $M$ can be used to model a physical process in the laboratory (e.g. the preparation of $\mathsf{A}$ conditioned on a classical variable), it does not correspond to a simple coarse-graining process applied to the classical degree of freedom.